# Comparing next-generation detector configurations for high-redshift gravitational wave sources with neural posterior estimation

### v1

@@ IV Conclusions
Given the unique sensitivity that ET and other XG gravitational wave detectors will provide to early-Universe stellar and primordial black holes, as well as to intermediate-mass black-hole binaries, we focus our analysis on BBH mergers with detector-frame chirp masses $\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$. Although this study addresses a more restricted science case than the full range of sources observable with XG detectors (Branchesi:2023mws; Abac:2025saz), it targets a regime with strong potential for breakthrough discoveries.
We assess the performance of seven configurations of XG detectors using neural posterior estimation enhanced with importance sampling, as implemented in Dingo-IS (Dax:2022pxd).
@@
We conduct a large injection campaign by sampling 1000 BBH merger parameters from the prior. As shown in Figure 1, Dingo-IS achieves excellent performance, recovering 85–96% of signals with a sample efficiency above 1%, depending on the detector configuration and for median SNRs between 40 and 70. This demonstrates that its efficiency remains high even for high-SNR events.
In Section III.2, we present a single-injection study showing that Dingo-IS accurately reproduces complex, disconnected posteriors consistent with standard stochastic-sampling methods, but at a fraction of the computational cost—few minutes instead of several hours.
@@
Our analysis shows that a network of two misaligned L-shaped ET detectors (2L MisA) can produce multimodal luminosity–distance posteriors (see top panel of Figure 2), and its distance estimates are generally less precise than those obtained with the triangular ET design ($\Delta$). However, Figure 3 demonstrates that the 2L MisA outperforms the $\Delta$ configuration in terms of information gain, sky localization, and volume reconstruction. The $\Delta$ configuration’s larger number of sky modes leads to poorer sky-localization performance, which in turn dominates its volume uncertainty.
@@
For high-mass and high-redshift BBH events, the fraction of sources recovered within a single sky mode remains low for all ET configurations ($\Delta$, 2L A, and 2L MisA). It surpasses 50% when CE is added to the global XG network and exceeds 70% when, instead, three LIGO detectors operate at A# sensitivity, as shown in Figure 4.
@@
We thank Stephen R. Green, Francesco Iacovelli, Luca Reali, Alexandre Toubiana, Manuel Arca Sedda, Michele Maggiore, Archisman Gosh, Elisa Maggio and Emanuele Berti for useful discussions. F.S. has been funded by the European Union – NextGenerationEU under the Italian Ministry of University and Research (MUR) ”Decreto per l’assunzione di ricercatori internazionali post-dottorato PNRR” - Missione 4 ”Istruzione e Ricerca” Componente 2 ”Dalla Ricerca all’Impresa” del PNRR - Investimento 1.2 “Finanziamento di progetti presentati da giovani ricercatori” - CUP D13C25000700001. M.B. and J.H. acknowledge support from the Astrophysics Center for Multi-messenger Studies in Europe (ACME), funded under the European Union’s Horizon Europe Research and Innovation Program, Grant Agreement No. 101131928. The data analysis has been performed on the HPC workstations ”Merac2A” financially supported by the MERAC Foundation and the European Astronomical Society through the 2023 MERAC Prize and the MERAC Support to research (PI Manuel Arca Sedda). The research leading to these results has been conceived and developed within the Einstein Telescope Observational Science Board ET-0611A-25.
@@ Appendix A Detector and network configurations
We provide further details on the detector and network configurations considered in this study. We limit the frequency range from $f_{\mathrm{min}}=6$ Hz to $f_{\mathrm{max}}=256$ Hz, using 8-second time segments, which results in frequency bins of size $\mathrm{d}f=0.125$ Hz. Setting the lower frequency limit to $f_{\mathrm{min}}=6$ Hz restricts the duration of all processed signals to less than 8 seconds. This approximation slightly underestimates ET performance, given its expected sensitivity down to $\sim 2$ Hz, but remains adequate for CE, whose sensitivity declines markedly below 5 Hz.
All ET configurations ($\Delta$, 2L A, and 2L MisA) adopt the HFLF cryogenic ASD curve (et), which incorporates both a high-frequency (HF) instrument and a cryogenic low-frequency (LF) instrument. Locations and orientations for the ET configurations are taken from Branchesi:2023mws. The ASDs for the three LIGO observatories (Livingston, Hanford, and India) at A# sensitivity is taken from ligo-india-asd, while the coordinates and orientation of the Indian interferometer from ligo-india. The location of CE is assumed to be the near the LIGO-Hanford Observatory (LHO), while the ASD for the 40-km arms in the baseline configuration is taken from ce_asd. Figure 5 shows the amplitude spectral densities (ASDs) used in this study for each detector over the specified frequency range.
@@
Figure 5: Amplitude spectral densities (ASDs) adopted in this work for different XG detectors. See Section II.1 and Appendix A for details.
@@
Table 2: Locations (latitude and longitude), orientations ($x$-arm azimuth), and chord distances for the different configurations considered in this work. See Section II.1 and Appendix A for details. Notes: The $x$-arm azimuth is measured counterclockwise from East. For all detectors, the $y$-arm azimuth is given by $x$-arm azimuth + 90°, except for the $\Delta$ configuration, where the $y$-arm azimuth = $x$-arm azimuth + 60°.
@@ Appendix B Waveforms and priors
We define spin-aligned, quasi-circular waveforms defined by the parameter set $\theta=\{\mathcal{M}_{\mathrm{d}}$, $q$, $D_{\mathrm{L}}$, $\mathrm{dec}$, $\mathrm{ra}$, $\theta_{\mathrm{jn}}$, $\psi$, $t_{\mathrm{geocent}}$, $\chi_{1}$, $\chi_{2}$, $\phi_{c}\}$, where $\mathcal{M}_{\mathrm{d}}$, $q$, and $D_{\mathrm{L}}$ denote the detector-frame chirp mass, mass ratio, and luminosity distance, respectively; $\mathrm{ra}$ and $\mathrm{dec}$ represent the right ascension and declination of the source; $\theta_{\mathrm{jn}}$ denotes the angle between the line of sight and the total angular momentum $\vec{J}$. In the absence of precession, as assumed in this work, $\vec{J}$ is equal to the orbital angular momentum of the binary $\vec{L}$, and thus $\theta_{\mathrm{jn}}$ corresponds to the inclination angle $\iota$; $\psi$ is the polarization angle; $t_{\mathrm{geocent}}$ is the coalescence time at the Earth’s center; and $\phi_{c}$ is the phase at coalescence. The waveforms also depend on the aligned-spin parameters $\chi_{1}$ and $\chi_{2}$, defined as $\chi_{i}=a_{i}\cos\theta_{i}$, where $a_{i}=|\chi_{i}|$ is the spin magnitude and $\theta_{i}$ is the angle with respect to the orbital angular momentum.
@@
We employ frequency-domain waveforms with a reference frequency fixed at $f_{\mathrm{ref}}=20$ Hz (Varma:2021csh), using the IMRPhenomXPHM waveform approximant (PhysRevD.103.104056), which includes the subdominant modes $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4)]$ in addition to the dominant $(\ell,|m|)=(2,2)]$ mode.
The prior distributions $\pi(\theta)$ employed in training Dingo are reported in Table 3. In particular, for the detector-frame chirp mass and mass ratio, we choose a prior uniform in primary and secondary masses (GWFish_Priors_Tutorial), therefore:
$\pi(\mathcal{M}_{\mathrm{d}})\propto\mathcal{M}_{\mathrm{d}},$ (B1)
where $\mathcal{M}_{\mathrm{d}}\in[40,~1100]$ M; and:
$\pi(q)\propto\frac{(1+q)^{2/5}}{q^{6/5}},$ (B2)
@@
where $q\in[0.125,~1]$. For the luminosity distance, we choose a prior uniform in comoving volume and source-frame time,
$\pi(D_{\mathrm{L}})\propto\frac{\mathrm{d}V_{c}}{\mathrm{d}D_{\mathrm{L}}}\frac{1}{1+z},$ (B3)
@@
where $D_{\mathrm{L}}\in[5,~500]$ Gpc, assuming cosmological parameters as provided by Planck:2018vyg. For the aligned spin component $\chi_{i}$, we form a joint prior as follows:
$\pi(\chi_{i})=\int_{0}^{0.9}\mathrm{d}a_{i}\int_{-1}^{1}\mathrm{d}\cos\theta_{i}\pi(a_{i})\pi(\cos\theta_{i})\delta(\chi_{i}-a_{i}\cos\theta_{i}),$ (B4)
where $\pi(a_{i})=\mathcal{U}(0,0.9)$ and $\cos\theta_{i}=\mathcal{U}(-1,1)$ (lange2018rapidaccurateparameterinference; Ashton:2018jfp).
@@
By convention, we set the Earth’s orientation and the positions and orientations of interferometers to those corresponding to the reference time $t_{\mathrm{ref}}=1126259462.391$ s, which coincides with the GPS trigger time of GW150914 (Romero-Shaw:2020owr). The merger time ($t_{\mathrm{geocent}}$) of each event is randomly sampled around this reference time. Fixing $t_{\mathrm{ref}}$ does not entail any loss of generality for short signals, as any time offset can be recovered.
@@
$\mathcal{M}_{\mathrm{d}}$ M Eq. B1
$q$ Eq. B2
$D_{\mathrm{L}}$ Gpc Eq. B3
@@
$\chi_{i}$ Eq. B4
Table 3: Parameters defining GW signals and their adopted priors. $\mathcal{U}(a,b)$ represents a uniform distribution between $a$ and $b$. See Appendix B for details.
@@ Appendix C Learning curves and p-p plots
Figure 6 shows that training and test losses remain consistent across all considered detector configurations, indicating the absence of overfitting. We observe an improvement in the training process compared to Figure 3 of Santoliquido:2025lot, where the test losses exhibited significant noise before epoch $\sim 200$.
Figure 7 displays the probability–probability (P–P) plots for the different detector configurations. P–P plots are commonly used to assess the reliability of inference algorithms (Cook01092006; Veitch:2014wba; talts2020validatingbayesianinferencealgorithms; 2020PhRvD.102j4057G), by verifying whether the fraction of injected parameters recovered within a given credible interval follows a uniform distribution. This method also enables the computation of $p$-values for each parameter through a Kolmogorov–Smirnov test. The results reported in Table 4 show good overall consistency. The plots in Figure 7 demonstrate that the improved training procedure produces an unbiased proposal distribution $q_{\phi}(\theta|d)$, as evidenced by the curves lying close to a uniform distribution.
Figure 6: The log loss as a function of training epochs shown for all detector configurations considered in this study, with the solid line representing the training set and the dashed line representing the test set. See Appendix C for further details.
@@
Table 4: $p$-values (color-coded) for each GW parameter (rows) across different detector configurations (columns). See Figure 7 and Appendix C for details.
Appendix D Dingo-IS vs. Bilby
@@
Table 5: Injected parameters $\theta^{\mathrm{inj}}$ of the event shown in Figure 2, along with the median Jensen–Shannon divergence ($\langle\mathrm{JSD}\rangle$, in units of $10^{-4}$ nat), which quantifies the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. For reference, the corresponding JSD threshold (JSDthr, in units of $10^{-4}$ nat) is also shown. The median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS are reported for two different detector configurations: 2L MisA and 2L MisA + CE. The bottom rows display the optimal SNR ($\rho$, Equation D4), source-frame primary ($m_{1,\mathrm{s}}$) and secondary ($m_{2,\mathrm{s}}$) masses, redshift ($z$), sky and comoving volume localization error ($\Delta\Omega_{90\%}$ and $\Delta V^{c}_{90\%}$), sample efficiency (Equation 2), log evidence and its uncertainty computed with Dingo-IS (Equation D5) and Bilby, and computing time in Bilby obtained using npool = 8 parallel processes. See Appendix D for details.
Figure 8: Marginalized one- and two-dimensional posterior distributions for all parameters for the event shown in Figure 2 observed with the 2L MisA configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
Figure 9: Marginalized one- and two-dimensional posterior distributions for all parameters of the same event shown in Figure 8, this time observed with the 2L MisA + CE configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
The parameters $\theta$ of a GW signal are inferred from strain data $d$ through Bayes’ theorem (vanderSluys:2007st; vanderSluys:2009bf; Thrane:2018qnx; Christensen:2022bxb):
$p(\theta|d)=\frac{\mathcal{L}(d|\theta)\pi(\theta)}{\mathcal{Z}(d)},$ (D1)
where $\mathcal{L}(d|\theta)$ is the likelihood function, $\pi(\theta)$ is the prior distribution (see Appendix B) and $\mathcal{Z}(d)=\int\mathrm{d}\theta\mathcal{L}(d|\theta)\pi(\theta)$ is the evidence.
Assuming stationary Gaussian noise, the likelihood in Equation D1 takes the form (Finn:1992wt; Veitch:2014wba; Romano:2016dpx; Ashton:2018jfp):
$\mathcal{L}(d|\theta)\propto\exp\left[\sum_{i}\left((d_{i}|h_{i}(\theta))-\frac{1}{2}(h_{i}(\theta)\,|\,h_{i}(\theta))\right)\right],$ (D2)
@@
$(a|b)=4\,\mathrm{Re}\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\mathrm{d}f\,\frac{a^{*}(f)\,b(f)}{S(f)},$ (D3)
@@
with $\mathrm{Re}$ indicating the real part and the complex conjugate. The quantity $\sqrt{S(f)}$ represents the noise ASD (see Appendix A for details). Equation D3 is also used to compute the optimal signal-to-noise ratio (SNR):
$\rho=\sqrt{(h(\theta^{\mathrm{inj}})|h(\theta^{\mathrm{inj}}))},$ (D4)
@@
where $h(\theta^{\mathrm{inj}})$ denotes the GW signal evaluated with injected parameters. For a network of detectors, the total optimal SNR is obtained by adding the individual detector SNRs in quadrature.
@@
To sample from Equation D1, we employ Bilby (Romero-Shaw:2020owr), a Python library designed for Bayesian inference with the nested sampling algorithm nessai (nessai). Thanks to its use of normalizing flows, nessai achieves rapid convergence. We fix the number of live points to n_live = 10_000.
@@
For the source shown in Figure 2 and observed with the 2L MisA configuration, the default nessai settings yield suboptimal performance. To improve robustness, we increase the volume_fraction from 0.95 to 0.98. This parameter controls the fraction of the total probability enclosed by the contour in the latent space—the base space from which samples are initially drawn before being mapped to physical space through the normalizing flow (Williams:2021qyt). Although a higher volume fraction improves performance, it also increases computational cost, which explains the long runtime ($\sim$ 50 hours) required for this source (see Table 5).
@@
To quantify the agreement between posterior samples obtained with Dingo-IS and Bilby, we employ the Jensen–Shannon divergence (JSD; 61115), a symmetrized variant of the Kullback–Leibler divergence (10.1214/aoms/1177729694). The JSD provides a measure of the similarity between two probability distributions, taking values from 0 nat, when the distributions are identical, up to $\ln(2)=0.69$ nat for completely distinct distributions.
@@
To determine a threshold for the JSD beyond which one-dimensional posteriors can be regarded as statistically different, we follow the procedure outlined in Santoliquido:2025lot. Using Dingo-IS, we generate 100 independent sets of $10^{5}$ posterior samples. For each parameter $\theta$, we compute the JSD across all 4950 unique pairs of these sets. Because all samples originate from the same underlying distribution, the resulting JSD values capture statistical fluctuations rather than meaningful discrepancies. We adopt the 95th percentile of these values as the threshold for each parameter (JSDthr in Table 5). Finally, we compare the one-dimensional posteriors obtained with Bilby to each Dingo-IS set, and verify whether the median JSD lies below the corresponding threshold ($\langle\mathrm{JSD}\rangle$ in Table 5).
@@
Table 5 lists the JSD values computed for the marginal one-dimensional posteriors of the source analyzed in Figure 2. The results show that the posteriors obtained with Bilby and Dingo-IS are statistically indistinguishable for both the 2L MisA and 2L MisA + CE configurations. This excellent agreement is further illustrated in Figures 8 and 9, which display the marginalized one- and two-dimensional posteriors for all parameters for the 2L MisA and 2L MisA + CE networks, respectively.

### v2

@@ 4 Conclusions
Given the unique sensitivity that ET and other XG GW detectors will provide for early-Universe stellar and PBHs, as well as IMBH binaries, we focused our analysis on BBH mergers with detector-frame chirp masses ($\mathcal{M}_{\mathrm{d}}$) $100\penalty 10000\ \mathrm{M}_{\odot}$. Although this study addresses a more restricted science case than the full range of sources observable with XG detectors (Branchesi and others, 2023; Abac and others, 2025c), it targets a regime with strong potential for breakthrough discoveries.
@@
We assessed the performance of seven configurations of XG detectors using NPE enhanced with importance sampling, as implemented in Dingo-IS (Dax et al., 2023). We conduced a large injection campaign by sampling 1000 BBH merger parameters from the prior. As shown in Fig. 1, Dingo-IS achieves excellent performance, recovering 85–96% of signals with a sample efficiency above 1%, depending on the detector configuration, for median S/Ns between 40 and 70. This demonstrates that its efficiency remains high even for high-S/N events.
@@
In Sect. 3.2 we present a single-injection study showing that Dingo-IS accurately reproduces complex, disconnected posteriors consistent with standard stochastic-sampling methods but at a fraction of the computational cost—a few minutes instead of several hours. Our analysis shows that a network of two misaligned L-shaped ET detectors (2L MisA) can produce multimodal luminosity–distance posteriors (see the top panel of Fig. 2), and its distance estimates are generally less precise than those obtained with the triangular ET design ($\Delta$). However, Fig. 3 demonstrates that the 2L MisA outperforms the $\Delta$ configuration in terms of information gain, sky localization, and volume reconstruction. The $\Delta$ configuration’s larger number of sky modes leads to poorer sky-localization performance, which in turn dominates its volume uncertainty.
@@
For both high-mass and high-redshift BBH events, the fraction of sources recovered within a single sky mode remains low for all ET configurations ($\Delta$, 2L A, and 2L MisA). It surpasses 50% when CE is added to the global XG network and exceeds 70% when, instead, three LIGO detectors operate at A# sensitivity, as shown in Fig. 4.
Data availability
The data used in this study are publicly available on Zenodo. The dataset corresponding to the $\Delta$ configuration can be found at Santoliquido (2026a, https://zenodo.org/records/18911700); the datasets for the 2L A, 2L MisA, and $\Delta$ + CE configurations are available at Santoliquido (2026b, https://zenodo.org/records/18911495); and the datasets for 2L MisA + LHI, 1L + CE, and 2L MisA + CE are available at Santoliquido (2026c, https://zenodo.org/records/18911380). The latest public release of Dingo is available at https://github.com/dingo-gw/dingo; this work is based on commit df6e586 in the fork Dingo-ET (https://github.com/filippo-santoliquido/dingo-ET). Additional data and code are available from the corresponding authors upon reasonable request.
Acknowledgements.
@@
We thank the anonymous referee for their thorough and careful review of our manuscript. We thank Stephen R. Green, Francesco Iacovelli, Luca Reali, Alexandre Toubiana, Manuel Arca Sedda, Michele Maggiore, Archisman Gosh, Elisa Maggio and Emanuele Berti for useful discussions. F.S. has been funded by the European Union – NextGenerationEU under the Italian Ministry of University and Research (MUR) ”Decreto per l’assunzione di ricercatori internazionali post-dottorato PNRR” - Missione 4 ”Istruzione e Ricerca” Componente 2 ”Dalla Ricerca all’Impresa” del PNRR - Investimento 1.2 “Finanziamento di progetti presentati da giovani ricercatori” - CUP D13C25000700001. M.B. and J.H. acknowledge support from the Astrophysics Center for Multi-messenger Studies in Europe (ACME), funded under the European Union’s Horizon Europe Research and Innovation Program, Grant Agreement No. 101131928. The data analysis has been performed on the HPC workstations ”Merac2A” financially supported by the MERAC Foundation and the European Astronomical Society through the 2023 MERAC Prize and the MERAC Support to research (PI Manuel Arca Sedda). The research leading to these results has been conceived and developed within the Einstein Telescope Observational Science Board ET-0611A-25.
@@ Appendix A Detector and network configurations
We provide further details on the detector and network configurations considered in this study. We limited the frequency range from $f_{\mathrm{min}}=6$ Hz to $f_{\mathrm{max}}=256$ Hz, using 8-second time segments, which results in frequency bins of size $\mathrm{d}f=0.125$ Hz. Setting the lower frequency limit to $f_{\mathrm{min}}=6$ Hz restricts the duration of all processed signals to less than 8 seconds. This approximation slightly underestimates ET performance, given its expected sensitivity down to $\sim 2$ Hz, but remains adequate for CE, whose sensitivity declines markedly below 5 Hz.
All ET configurations ($\Delta$, 2L A, and 2L MisA) adopt the high frequency–low frequency (HFLF) cryogenic ASD curve (ET Collaboration, 2023), which incorporates both a high-frequency instrument and a cryogenic low-frequency instrument. Locations and orientations for the ET configurations are taken from Branchesi and others (2023). The ASDs for the three LIGO observatories (Livingston, Hanford, and India) at A# sensitivity is taken from LIGO Collaboration (2020), while the coordinates and orientation of the Indian interferometer from Kandhasamy and Bose (2020). The location of CE is assumed to be the near the LIGO-Hanford Observatory (LHO), while the ASD for the 40-km arms in the baseline configuration is taken from CE Collaboration (2022). Figure 6 shows the ASDs used in this study for each detector over the specified frequency range.
@@
Figure 6: ASDs adopted in this work for different XG detectors. See Sect. 2.1 and Appendix A for details.
Table 2: Locations (latitude and longitude), orientations ($x$-arm azimuth), and chord distances for the different detector configurations considered in this work. [^2] The $x$-arm azimuth is measured counterclockwise from East. For all detectors, the $y$-arm azimuth is given by $x$-arm azimuth + 90°, except for the $\Delta$ configuration, where the $y$-arm azimuth = $x$-arm azimuth + 60°. See Sect. 2.1 and Appendix A for details.
@@ Appendix B Waveforms and priors
We define spin-aligned, quasi-circular waveforms defined by the parameter set $\theta=\{\mathcal{M}_{\mathrm{d}}$, $q$, $D_{\mathrm{L}}$, $\mathrm{dec}$, $\mathrm{ra}$, $\theta_{\mathrm{jn}}$, $\psi$, $t_{\mathrm{geocent}}$, $\chi_{1}$, $\chi_{2}$, $\phi_{c}\}$, where $\mathcal{M}_{\mathrm{d}}$, $q$, and $D_{\mathrm{L}}$ denote the detector-frame chirp mass, mass ratio, and luminosity distance, respectively; $\mathrm{ra}$ and $\mathrm{dec}$ represent the right ascension and declination of the source; $\theta_{\mathrm{jn}}$ denotes the angle between the line of sight and the total angular momentum (J\@vec{J}). In the absence of precession, as assumed in this work, J\@vec{J} is equal to the orbital angular momentum of the binary (L\@vec{L}), and thus $\theta_{\mathrm{jn}}$ corresponds to the inclination angle ($\iota$); $\psi$ is the polarization angle; $t_{\mathrm{geocent}}$ is the coalescence time at the Earth’s center; and $\phi_{c}$ is the phase at coalescence. The waveforms also depend on the aligned-spin parameters $\chi_{1}$ and $\chi_{2}$, defined as $\chi_{i}=a_{i}\cos\theta_{i}$, where $a_{i}=|\chi_{i}|$ is the spin magnitude and $\theta_{i}$ is the angle with respect to the orbital angular momentum.
@@
We employ frequency-domain waveforms with a reference frequency fixed at $f_{\mathrm{ref}}=20$ Hz (Varma et al., 2022), using the IMRPhenomXPHM waveform approximant (Pratten et al., 2021; Colleoni et al., 2025), which includes the subdominant modes $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4)]$ in addition to the dominant $(\ell,|m|)=(2,2)]$ mode. The prior distributions, $\pi(\theta),$ employed in training Dingo are reported in Table 3. In particular, for the detector-frame chirp mass and mass ratio, we chose a prior uniform in primary and secondary masses (Dupletsa, 2025); therefore,
$\pi(\mathcal{M}_{\mathrm{d}})\propto\mathcal{M}_{\mathrm{d}},$ (4)
where $\mathcal{M}_{\mathrm{d}}\in[40,\penalty 10000\ 1100]$ M; and
$\pi(q)\propto\frac{(1+q)^{2/5}}{q^{6/5}},$ (5)
@@
where $q\in[0.125,\penalty 10000\ 1]$. For the luminosity distance, we chose a prior uniform in comoving volume and source-frame time,
$\pi(D_{\mathrm{L}})\propto\frac{\mathrm{d}V_{c}}{\mathrm{d}D_{\mathrm{L}}}\frac{1}{1+z},$ (6)
@@
where $D_{\mathrm{L}}\in[5,\penalty 10000\ 500]$ Gpc, assuming cosmological parameters as provided by Aghanim and others (2020). For the aligned spin component $\chi_{i}$, we form a joint prior as follows:
$\pi(\chi_{i})=\int_{0}^{0.9}\mathrm{d}a_{i}\int_{-1}^{1}\mathrm{d}\cos\theta_{i}\pi(a_{i})\pi(\cos\theta_{i})\delta(\chi_{i}-a_{i}\cos\theta_{i}),$ (7)
where $\pi(a_{i})=\mathcal{U}(0,0.9)$ and $\cos\theta_{i}=\mathcal{U}(-1,1)$ (Lange et al., 2018; Ashton and others, 2019).
@@
By convention, we set the Earth’s orientation and the positions and orientations of interferometers to those corresponding to the reference time $t_{\mathrm{ref}}=1126259462.391$ s, which coincides with the GPS trigger time of GW150914 (Romero-Shaw and others, 2020). The merger time ($t_{\mathrm{geocent}}$) of each event was randomly sampled around this reference time. Fixing $t_{\mathrm{ref}}$ does not entail any loss of generality for short signals, as any time offset can be recovered.
[…]
@@
$\mathcal{M}_{\mathrm{d}}$ M Eq. 4
$q$ Eq. 5
$D_{\mathrm{L}}$ Gpc Eq. 6
@@
$\chi_{i}$ Eq. 7
@@ Appendix C Learning curves and P-P plots
Figure 7 shows that training and test losses remain consistent across all considered detector configurations, indicating the absence of overfitting. We observe an improvement in the training process compared to Fig. 3 of Santoliquido and others (2025), where the test losses exhibited significant noise before epoch $\sim 200$.
Figure 8 displays the probability–probability (P–P) plots for the different detector configurations. P–P plots are commonly used to assess the reliability of inference algorithms (Samantha R Cook and Rubin, 2006; Veitch and others, 2015; Talts et al., 2020; Green et al., 2020a), by verifying whether the fraction of injected parameters recovered within a given credible interval follows a uniform distribution. This method also enables the computation of $p$-values for each parameter through a Kolmogorov–Smirnov test. The results reported in Table 4 show good overall consistency. The plots in Fig. 8 demonstrate that the improved training procedure produces an unbiased proposal distribution $q_{\phi}(\theta|d)$, as evidenced by the curves lying close to a uniform distribution.
Figure 7: Log loss as a function of training epochs shown for all detector configurations considered in this study. The solid line represents the training set and the dashed line the test set. See Appendix C for further details.
@@
Table 4: $p$-values (color-coded) for each GW parameter (rows) across different detector configurations (columns). [^4] See Fig. 8 and Appendix C for details.
@@
Appendix D Dingo-IS versus Bilby
Table 5: Injected parameters ($\theta^{\mathrm{inj}}$) of the event shown in Fig. 2, along with the median JSD ($\langle\mathrm{JSD}\rangle$, in units of $10^{-4}$ nat), which quantifies the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. [^5] For reference, the corresponding JSD threshold (JSDthr, in units of $10^{-4}$ nat) is also shown. The median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS are reported for two different detector configurations: 2L MisA and 2L MisA + CE. The bottom rows display the optimal S/N ($\rho$, Eq. 11), source-frame primary ($m_{1,\mathrm{s}}$) and secondary ($m_{2,\mathrm{s}}$) masses, redshift ($z$), sky and comoving volume localization error ($\Delta\Omega_{90\%}$ and $\Delta V^{c}_{90\%}$), sample efficiency (Eq. 2), log evidence and its uncertainty computed with Dingo-IS (Eq. 12) and Bilby, and computing time in Bilby obtained using npool = 8 parallel processes. See Appendix D for details.
@@ Appendix D Dingo-IS versus Bilby
Figure 9: Marginalized one- and two-dimensional posterior distributions for all parameters for the event shown in Fig. 2 observed with the 2L MisA configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
Figure 10: Marginalized one- and two-dimensional posterior distributions for all parameters of the same event shown in Fig. 9, this time observed with the 2L MisA + CE configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
The parameters ($\theta$) of a GW signal are inferred from strain data ($d$) through Bayes’ theorem (van der Sluys et al., 2008, 2009; Thrane and Talbot, 2019; Christensen and Meyer, 2022):
$p(\theta|d)=\frac{\mathcal{L}(d|\theta)\pi(\theta)}{\mathcal{Z}(d)},$ (8)
where $\mathcal{L}(d|\theta)$ is the likelihood function, $\pi(\theta)$ is the prior distribution (see Appendix B) and $\mathcal{Z}(d)=\int\mathrm{d}\theta\mathcal{L}(d|\theta)\pi(\theta)$ is the evidence. Assuming stationary Gaussian noise, the likelihood in Eq. 8 takes the form (Finn, 1992; Veitch and others, 2015; Romano and Cornish, 2017; Ashton and others, 2019)
$\mathcal{L}(d|\theta)\propto\exp\left[\sum_{i}\left((d_{i}|h_{i}(\theta))-\frac{1}{2}(h_{i}(\theta)\,|\,h_{i}(\theta))\right)\right],$ (9)
@@
$(a|b)=4\,\mathrm{Re}\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\mathrm{d}f\,\frac{a^{*}(f)\,b(f)}{S(f)},$ (10)
@@
with $\mathrm{Re}$ indicating the real part and the complex conjugate. The quantity $\sqrt{S(f)}$ represents the noise ASD (see Appendix A for details). Equation 10 is also used to compute the optimal S/N:
$\rho=\sqrt{(h(\theta^{\mathrm{inj}})|h(\theta^{\mathrm{inj}}))},$ (11)
@@
where $h(\theta^{\mathrm{inj}})$ denotes the GW signal evaluated with injected parameters. For a network of detectors, the total optimal S/N is obtained by adding the individual detector S/Ns in quadrature.
@@
To sample from Eq. 8, we employed Bilby (Romero-Shaw and others, 2020), a Python library designed for Bayesian inference with the nested sampling algorithm nessai (Williams, 2021). Thanks to its use of normalizing flows, nessai achieves rapid convergence. We fixed the number of live points to n_live = 10_000.
@@
For the source shown in Fig. 2 and observed with the 2L MisA configuration, the default nessai settings yield suboptimal performance. To improve robustness, we increased the volume_fraction from 0.95 to 0.98. This parameter controls the fraction of the total probability enclosed by the contour in the latent space—the base space from which samples are initially drawn before being mapped to physical space through the normalizing flow (Williams et al., 2021). Although a higher volume fraction improves performance, it also increases computational cost, which explains the long runtime ($\sim$ 50 hours) required for this source (see Table 5).
@@
To quantify the agreement between posterior samples obtained with Dingo-IS and Bilby, we employ the Jensen–Shannon divergence (JSD; Lin, 1991), a symmetrized variant of the KL divergence (Kullback and Leibler, 1951). The JSD provides a measure of the similarity between two probability distributions, taking values from 0 nat, when the distributions are identical, up to $\ln(2)=0.69$ nat for completely distinct distributions.
@@
To determine a threshold for the JSD beyond which one-dimensional posteriors can be regarded as statistically different, we follow the procedure outlined in Santoliquido and others (2025). Using Dingo-IS, we generate 100 independent sets of $10^{5}$ posterior samples. For each parameter $\theta$, we compute the JSD across all 4950 unique pairs of these sets. Because all samples originate from the same underlying distribution, the resulting JSD values capture statistical fluctuations rather than meaningful discrepancies. We adopt the 95th percentile of these values as the threshold for each parameter (JSDthr in Table 5). Finally, we compare the one-dimensional posteriors obtained with Bilby to each Dingo-IS set, and verify whether the median JSD lies below the corresponding threshold ($\langle\mathrm{JSD}\rangle$ in Table 5).
@@
Table 5 lists the JSD values computed for the marginal one-dimensional posteriors of the source analyzed in Fig. 2. The results show that the posteriors obtained with Bilby and Dingo-IS are statistically indistinguishable for both the 2L MisA and 2L MisA + CE configurations. This excellent agreement is further illustrated in Figs. 9 and 10, which display the marginalized one- and two-dimensional posteriors for all parameters for the 2L MisA and 2L MisA + CE networks, respectively.